Proceedings of PASCO’94 \begin{matrix}c \\ ◯ \end{matrix} 1994 by World Scientific Publishing Company

The Title

First Author’s Name First Author’s Address Second Author’s Name Second Author’s Address

Abstract: The Abstract.

Keywords: The Key Words

1 Introduction

The Introduction. As described in the CFP:

The introduction must consist of
a brief and intuitive description of:
* the problem that is tackled in the
  paper
* the motivation and importance
* the method used in the paper
* the performance or the result
* the related works and the
  originality claim

2 First Main Section

blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah

blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah

blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah

3 Second Main Section

blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah

blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah

blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah …

References

[1] E. Xample. How to write good example files. In Proceedings of PASCO’94, Hagenberg, Austria, 1994.